OpenStudy (lisa123):

what will be the amount of sugar in milligrams if the size of the milk chocolate bar is reduced from 17.00g to 5.00g

OpenStudy (calculusxy):

First we should convert the grams into mg... 1 gram = 1000 mg 17 g = 17,000mg 5 g = 5,000mg

OpenStudy (calculusxy):

And then you can subtract 17,000mg - 5,000mg

OpenStudy (calculusxy):

did u get an answer?

OpenStudy (calculusxy):

@lisa123

OpenStudy (lisa123):

12,000mg??

OpenStudy (calculusxy):

yes!

OpenStudy (calculusxy):

oh wait....

OpenStudy (calculusxy):

if it is just saying to know in mg after it got reduced to 5g, then it is only 5000mg

OpenStudy (lisa123):

no thats not the right answer

OpenStudy (lisa123):

originally there is 17g of milk and 10.5g of sugar in the milk

OpenStudy (phi):

do they tell how much sugar is in the larger bar ?

OpenStudy (calculusxy):

then i think that you can make a proportion

OpenStudy (lisa123):

thats what i thought idk how to do it

OpenStudy (lisa123):

@calculusxy @phi

OpenStudy (phi):

\[ \frac{5}{17} = \frac{x}{10500} \]

OpenStudy (lisa123):

17x/5(10500)??

OpenStudy (lisa123):

x=2.948

OpenStudy (lisa123):

294.1

OpenStudy (phi):

I would multiply both sides by 10500 \[ \frac{5}{17}\cdot 10500 = \frac{x}{10500}\cdot 10500 \]

OpenStudy (phi):

on the right side , 10500/10500 simplifies to 1 and 1*x is x the left side will be the answer in milligrams

OpenStudy (lisa123):

Thank you!

OpenStudy (phi):

what did you get ?

OpenStudy (lisa123):

3088mg

OpenStudy (phi):

yes, that looks good.